# Minimizing False-Positive Attributions in Explanations of Non-Linear Models

### v1

@@ Abstract
Suppressor variables can influence model predictions without being dependent on the target outcome and they pose a significant challenge for Explainable AI (XAI) methods. These variables may cause false-positive feature attributions, undermining the utility of explanations. Although effective remedies exist for linear models, their extension to non-linear models and to instance-based explanations has remained limited. We introduce PatternLocal, a novel XAI technique that addresses this gap. PatternLocal begins with a locally linear surrogate, e.g. LIME, KernelSHAP, or gradient-based methods, and transforms the resulting discriminative model weights into a generative representation, thereby suppressing the influence of suppressor variables while preserving local fidelity. In extensive hyperparameter optimization on the XAI-TRIS benchmark, PatternLocal consistently outperformed other XAI methods and reduced false-positive attributions when explaining non-linear tasks, thereby enabling more reliable and actionable insights.
@@ 1 Introduction
In recent years, substantial advancements in explainable artificial intelligence (XAI) have been driven by the growing demand for transparent, accountable, and ethical AI systems across various industries and research communities. The need to understand machine learning models’ decision-making processes is now well-established, particularly in high-stakes domains such as healthcare, finance, and criminal justice. While promising progress has been made with so-called self-explainable models such as prototypical networks, these approaches often fall short of achieving the same downstream performance as conventional models and typically require complete retraining [18]. Consequently, significant attention remains focused on post-hoc XAI methods that can be applied to existing models without retraining.
@@
Figure 1: Conceptual overview of how local linearization (top row) and PatternLocal (both rows) generate explanations. From the dataset (a), we train a classifier, whose decision boundaries are visualized using 2D projections. Traditional local XAI methods (top) approximate the model with a linear surrogate around the instance of interest. In contrast, PatternLocal (both rows) enhances this process by transforming the local discriminative surrogate into a generative explanation, effectively suppressing the impact of irrelevant or misleading features. The resulting importance maps (c, d) can be directly compared to the ground-truth attributions (b).
@@
Post-hoc XAI methods encompass a diverse array of approaches, including architecture-specific techniques such as Layer-wise Relevance Propagation (LRP) [1], concept-based explanations like TCAV [13], and model-agnostic frameworks such as SHAP [15] and LIME [17]. These methods promise to support model validation, data quality checks, and actionable intervention suggestions [19]. To provide evidence, XAI methods are evaluated against criteria such as faithfulness to the model, robustness to perturbations, and human interpretability [26]. However, none of these criteria guarantee that the features highlighted as important are discriminative for the target variable, an assumption behind many uses of XAI [12].
@@
Several studies [28, 29] have shown that widely used XAI methods often assign importance to suppressor variables [8, 27], features that affect model predictions without having a direct statistical dependency with the target. Suppressors may capture correlations in noise or other auxiliary signals that models exploit to improve predictions without offering causal or correlational insights. For instance, a model might predict epilepsy from seizure activity in a specific brain region but to optimally do so also rely on a noise probe from an unaffected region; common XAI methods may then highlight both regions, mistakenly assigning significance to the irrelevant area.
@@
To support interpretations that rely on statistical relevance, a feature should be dependent on the target variable [28, 12]. The Pattern method [11] formalizes this idea by separating prediction from explanation: a discriminative model makes predictions, and a corresponding generative model provides explanations. Extensions such as PatternNet and PatternAttribution generalize this idea to deep neural networks [14] but have shown weaknesses on non-linear suppressor benchmarks like XAI-TRIS [7]. Unlike human-centered evaluations, benchmarks like XAI-TRIS provide controlled environments with objective ground-truth explanations to evaluate the robustness of XAI methods against suppressors and other uninformative variables. Also, model-agnostic methods such as SHAP, LIME, and LRP largely ignore training data structure, making them prone to suppressor bias [28, 29, 7].
@@
In this work, we introduce PatternLocal, a data-driven, model-agnostic XAI method that extends the Pattern approach to non-linear cases. PatternLocal converts locally discriminative explanations, produced by methods such as LIME or KernelSHAP, into generative representations that more accurately reflect statistical relevance; see Figure 1. This transformation helps reduce the impact of suppressor and other non-informative variables while maintaining local fidelity to the model’s decision behavior. PatternLocal assumes access to representative training data and a meaningful input simplification (e.g. superpixels or feature masks) shared across samples. We evaluate PatternLocal on the XAI-TRIS suppressor and artificial lesion MRI benchmark and compare it with a range of established XAI methods. The results show that PatternLocal consistently provides more reliable and interpretable explanations without requiring any changes to the underlying model.
@@ 2.1 Notation
Let $\mathcal{X}\subset\mathbb{R}^{D}$ and $\mathcal{Y}\subset\mathbb{R}$ denote the input and output spaces, respectively. Consider a model $f:\mathcal{X}\rightarrow\mathcal{Y}$. For an instance $\mathbf{x_{\star}}\in\mathcal{X}$ we wish to explain the prediction $f(\mathbf{x_{\star}})$. Because highly complex models such as deep networks are notoriously hard to interpret, they are approximated locally by a simpler, intrinsically interpretable model $g$. Throughout we therefore restrict attention to local explanations for the specific instance $\mathbf{x_{\star}}$ rather than to the behavior of $f$ on $\mathcal{X}$ as a whole.
Many XAI methods first replace $\mathbf{x_{\star}}$ with a lower-dimensional or more semantically meaningful representation $\mathbf{x^{\prime}_{\star}}=h_{\mathbf{x_{\star}}}(\mathbf{x_{\star}})\in\mathbb{R}^{D^{\prime}}$. In image analysis, for example, $\mathbf{x^{\prime}_{\star}}$ may encode super-pixels; in other settings, it may correspond to a low-rank projection or simply a subset of raw features. Such simplification often makes for easier interpretation and, when $D^{\prime}\ll D$, also reduces the computational cost.
@@ 2.2 Suppressor Variables
Various studies [e.g., 8, 11] demonstrated that linear models may need to assign significant non-zero weight to correlated noise variables (suppressors). Given $\mathbf{x}\in\mathbb{R}^{D}$, assume each instance is generated from a linear model of the form
$\mathbf{x}\;=\;\mathbf{A}\,\mathbf{s}\;+\;\bm{\varepsilon},\qquad\mathbf{s}=\bigl{[}s_{1},\dots,s_{K}\bigr{]}^{\!\top}\in\mathbb{R}^{K},$ (1)
where $K$ is the number of latent factors, $\mathbf{A}=[\mathbf{a}_{1},\dots,\mathbf{a}_{K}]\in\mathbb{R}^{D\times K}$ contains the activation patterns $\mathbf{a}_{k}$, and $\bm{\varepsilon}$ denotes additive noise. A linear backward model attempts to recover the latent vector $\mathbf{s}$ via
$\hat{\mathbf{s}}\;=\;\mathbf{W}^{\!\top}\mathbf{x},\qquad\mathbf{W}\in\mathbb{R}^{D\times K}.$ (2)
Consider $D=2$ and $K=1$ with $x_{1}=s+d$ and $x_{2}=d$, where the scalar $s$ is the signal of interest and $d$ is a suppressor. The signal is exactly recovered by $\hat{s}=x_{1}-x_{2}=\mathbf{w}^{\!\top}\mathbf{x}$ with $\mathbf{w}=[1,-1]^{\!\top}$. Interpreting $\mathbf{w}$ directly as feature importance, however, suggests $x_{1}$ and $x_{2}$ contribute equally, even though the second channel carries only the suppressor, which is removed entirely in $\hat{s}$ by subtraction. This issue with suppressor variables extends to common explainable AI methods used to interpret machine learning models [29, 7].
@@
Instead of using the weights of the linear backward model, Haufe et al. [11] propose to use the corresponding forward model. They show that the weight matrix $\mathbf{W}$ has a unique forward-model counterpart
$\mathbf{A}\;=\;\Sigma_{\mathbf{X}}\,\mathbf{W}\,\Sigma_{\mathbf{S}}^{-1},$ (3)
where $\Sigma_{\mathbf{X}}=\operatorname{Cov}[\mathbf{x}]$ and $\Sigma_{\mathbf{S}}=\operatorname{Cov}[\hat{\mathbf{s}}]$ which recovers the activation pattern in Equation 1. If the estimated factors are uncorrelated, $\Sigma_{\mathbf{S}}$ is diagonal and Equation 3 reduces to $\mathbf{A}\;\propto\;\Sigma_{\mathbf{X}}\,\mathbf{W}\;=\;\operatorname{Cov}\bigl{[}\mathbf{x},\,\hat{\mathbf{s}}\bigr{]}$. The resulting global feature importance map for the entire data set $\mathcal{D}\subset\mathcal{X}$ is
@@ 2.4 Local explainability methods
$\mathbf{s}^{\text{LIME}}(\mathbf{x}_{\star})\;=\;\mathbf{w}^{\text{LIME}}\;=\;\arg\!\min_{\mathbf{v}}\;\mathbb{E}_{\mathbf{z}^{\prime}\,\sim\,\mathcal{Z}}\Bigl{[}\Pi_{\mathbf{x^{\prime}_{\star}}}\!\bigl{(}\mathbf{z}^{\prime}\bigr{)}\bigl{(}f\!\bigl{(}h_{\mathbf{x_{\star}}}^{-1}(\mathbf{z}^{\prime})\bigr{)}-\mathbf{v}^{\!\top}\mathbf{z}^{\prime}\bigr{)}^{2}\Bigr{]}\;+\;\lambda\,R(\mathbf{v}),$ (5)
where $R$ is a regularization weighted by $\lambda$, $\mathcal{Z}$ is the simplified input space, and $h_{\mathbf{x_{\star}}}^{-1}\!:\mathcal{Z}\!\rightarrow\!\mathcal{X}$ maps a simplified sample back to the original space. This notation is similar to the one by Tan et al. [25]. For image explanations the default choices are $\Pi_{\mathbf{x^{\prime}_{\star}}}(\mathbf{z}^{\prime})=\exp\,\!(-\|\mathbf{1}-\mathbf{z}^{\prime}\|_{0}^{2}/\sigma^{2})$, $R(\mathbf{v})=\|\mathbf{v}\|_{2}^{2}$, and $h_{\mathbf{x^{\prime}_{\star}}}^{-1}(\mathbf{z}^{\prime})=\mathbf{z}=\mathbf{x_{\star}}\odot B(\mathbf{z}^{\prime})+\mathbf{r}\odot(\mathbf{1}-B(\mathbf{z}^{\prime}))$, with $\mathbf{z}^{\prime}\sim\text{Uni}(\{0,1\}^{D^{\prime}})$, $\odot$ denoting element-wise product, $r$ being a reference value, and where $B:\{0,1\}^{d}\rightarrow\{0,1\}^{D}$ copies each bit $z^{\prime}_{j}$ to every pixel in super-pixel. As mentioned, LIME can assign non-zero importance to suppressor variables [29].
@@
As shown by Lundberg and Lee [15], Tan et al. [25], KernelSHAP can be written in the form of Equation 5 by setting $R(\mathbf{v})=0$ and $\Pi_{x^{\prime}}(z^{\prime})=(D^{\prime}-1)\big{/}\left[\left(D^{\prime}\,\mathrm{choose}\,\|z^{\prime}\|_{0}\right)\cdot\|z^{\prime}\|_{0}\cdot(D^{\prime}-\|z^{\prime}\|_{0})\right]$. Tan et al. [25] further demonstrate that Equation 5 unifies SmoothGrad [22] and the plain Gradient method [30].
@@ 3 PatternLocal
We now unify the Pattern insight that forward model parameters reveal statistically relevant features with the ability to probe highly non-linear models locally with methods like LIME. We name this PatternLocal. PatternLocal converts any linear surrogate explanation $\mathbf{w}$ produced around an instance $\mathbf{x}_{\star}$ into a data-driven pattern $\mathbf{a}$ that (i) suppresses false positives caused by suppressor variables, (ii) remains faithful to the model because it is built on the very surrogate that approximates the decision boundary, (iii) preserves local fidelity by weighting the surrounding data with a local kernel, and (iv) remains practical by operating in a simplified, low-dimensional space with regularized estimation.
(1) Suppressor-variable mitigation.
Pattern mitigates suppressor variables in linear models by converting the backward model into a forward one. This can be interpreted as regressing $\mathbf{x}$ on the estimated signal $\tilde{\mathbf{s}}=\mathbf{W}^{\top}\mathbf{x}$ under a Gaussian noise model. The goal is to reconstruct $\mathbf{x}$ from $\tilde{\mathbf{s}}$, effectively down-weighting features that do not carry the signal. PatternLocal extends this idea to the local setting by regressing the local surrogate output $\tilde{y}=\mathbf{w}^{\top}\mathbf{x}$ onto the neighborhood of $\mathbf{x}_{\star}$, thus removing suppressor effects specific to the instance.
(2) Model faithfulness.
PatternLocal inherits its surrogate $\mathbf{w}$ from Equation 5, which is explicitly optimized to approximate the model $f$ in the local neighborhood of $\mathbf{x}_{\star}$. By regressing the simplified input $\mathbf{x}$ on the surrogate model’s response $\tilde{y}=\mathbf{w}^{\top}\mathbf{x}$, PatternLocal preserves the local approximation. Crucially, we regress onto the surrogate output rather than the true label $y$, since $y$ reflects only the ground truth and not the behavior of the model $f$. Using $y$ instead would explain the task, not the model $f$, violating model faithfulness.
(3) Local fidelity.
In non-linear models, the important features can vary across different regions of the input space. PatternLocal enforces locality by weighting all terms in its regression objective using the local kernel $\Pi_{\mathbf{x^{\prime}_{\star}}}$. This ensures that the explanation focuses on the neighborhood around $\mathbf{x}_{\star}$ and does not reflect global structure.
(4) Feasibility and dimensionality.
PatternLocal requires access to the training data $\mathcal{D}$ and assumes that a sufficient number of samples lie in the neighborhood of $\mathbf{x}_{\star}$. To make this feasible, it operates in the same simplified space defined by $\mathbf{h}_{\mathbf{x_{\star}}}$ as in Equation 5, which not only aligns the explanation with the surrogate but also drastically reduces the dimensionality and the number of samples required. Nevertheless, the regression can still become ill-posed in sparse or high-dimensional settings. To ensure stability, we include a regularization $Q$.

### v3

@@ Abstract
Suppressor variables can influence model predictions without being dependent on the target outcome, and they pose a significant challenge for Explainable AI (XAI) methods. These variables may cause false-positive feature attributions, undermining the utility of explanations. Although effective remedies exist for linear models, their extension to non-linear models and instance-based explanations has remained limited. We introduce PatternLocal, a novel XAI technique that addresses this gap. PatternLocal begins with a locally linear surrogate, e.g., LIME, KernelSHAP, or gradient-based methods, and transforms the resulting discriminative model weights into a generative representation, thereby suppressing the influence of suppressor variables while preserving local fidelity. In extensive hyperparameter optimization on the XAI-TRIS benchmark, PatternLocal consistently outperformed other XAI methods and reduced false-positive attributions when explaining non-linear tasks, thereby enabling more reliable and actionable insights. We further evaluate PatternLocal on an EEG motor imagery dataset, demonstrating physiologically plausible explanations.
@@ 1 Introduction
In recent years, the growing demand for transparent, accountable, and ethical AI systems across various industries and research communities has driven substantial advancements in explainable artificial intelligence (XAI). The need to understand machine learning models’ decision-making processes is well-established, particularly in high-stakes domains such as healthcare, finance, and criminal justice. While researchers have made promising progress with so-called self-explainable models such as prototypical networks, these approaches often fail to achieve the same downstream performance as conventional models and typically require complete retraining [22]. Consequently, researchers focus significant attention on post-hoc XAI methods that apply to existing models without retraining.
@@
Figure 1: Conceptual step-by-step overview of how the PatternLocal method generates local explanations. A classifier is trained on the dataset (a), with both the dataset and the model’s decision boundary visualized in a 2D projection. Many XAI methods produce importance maps (c) through local linearization (e.g., LIME, KernelSHAP). PatternLocal enhances this process by transforming the local discriminative surrogate into a generative explanation (d), thereby reducing the influence of irrelevant or misleading features. The resulting importance maps (c–d) can be directly compared with the ground-truth attributions (b).
@@
Post-hoc XAI methods encompass a diverse array of approaches, including architecture-specific techniques such as Layer-wise Relevance Propagation (LRP) [1], concept-based explanations like TCAV [14], and model-agnostic frameworks such as SHAP [17] and LIME [21]. These methods promise to support model validation, data quality checks, and actionable intervention suggestions [23]. To provide evidence, researchers evaluate XAI methods against criteria such as faithfulness to the model, robustness to perturbations, and human interpretability [32]. However, none of these criteria guarantee that the features highlighted as important are discriminative for the target variable, an assumption behind many uses of XAI [13].
@@
Several studies [34, 35] have shown that widely used XAI methods often assign importance to suppressor variables [9, 33], features that affect model predictions without having a direct statistical dependency with the target. Suppressors may capture correlations in noise or other auxiliary signals that models exploit to improve predictions without offering causal or correlational insights. For instance, a model might predict epilepsy from seizure activity in a specific brain region, but to optimally do so, it also relies on a noise probe from an unaffected region; common XAI methods may then highlight both regions, mistakenly assigning significance to the irrelevant area.
@@
To support interpretations that rely on statistical relevance, a feature should depend on the target variable [34, 13]. The Pattern method [12] formalizes this idea by separating prediction from explanation: a discriminative model makes predictions, and a corresponding generative model provides explanations. Extensions such as PatternNet and PatternAttribution generalize this idea to deep neural networks [15] but have shown weaknesses on non-linear suppressor benchmarks like XAI-TRIS [8]. Unlike human-centered evaluations, benchmarks like XAI-TRIS provide controlled environments with objective ground-truth explanations to evaluate the robustness of XAI methods against suppressors and other uninformative variables. Also, model-agnostic methods such as SHAP, LIME, and LRP largely ignore training data structure, making them prone to suppressor bias [34, 35, 8].
@@
In this work, we introduce PatternLocal, a data-driven, model-agnostic XAI method that extends the Pattern approach to non-linear cases. PatternLocal converts locally discriminative explanations, produced by methods such as LIME or KernelSHAP, into generative representations that more accurately reflect statistical relevance; see Figure˜1. This transformation helps reduce the impact of suppressor and other non-informative variables while maintaining local fidelity to the model’s decision behavior. PatternLocal assumes access to representative training data and a meaningful input simplification (e.g., superpixels or feature masks) shared across samples. We evaluate PatternLocal on the XAI-TRIS benchmark, artificial lesion MRI benchmark, and an EEG Motor imagery dataset, and compare it with a range of established XAI methods. The results show that PatternLocal consistently provides more reliable and interpretable explanations without changing the underlying model.
@@ 2.1 Notation
Let $\mathcal{X}\subset\mathbb{R}^{D}$ and $\mathcal{Y}\subset\mathbb{R}$ denote the input and output spaces, respectively. Consider a model $f:\mathcal{X}\rightarrow\mathcal{Y}$. For an instance $\mathbf{x_{\star}}\in\mathcal{X}$ we wish to explain the prediction $f(\mathbf{x_{\star}})$. Because highly complex models such as deep networks are notoriously hard to interpret, they are approximated locally by a simpler, intrinsically interpretable model $g$. We therefore restrict attention to local explanations for the specific instance $\mathbf{x_{\star}}$ rather than to the behavior of $f$ on $\mathcal{X}$ as a whole.
Many XAI methods first replace $\mathbf{x_{\star}}$ with a lower-dimensional or more semantically meaningful representation $\mathbf{x^{\prime}_{\star}}=h_{\mathbf{x_{\star}}}(\mathbf{x_{\star}})\in\mathbb{R}^{D^{\prime}}$. In image analysis, for example, $\mathbf{x^{\prime}_{\star}}$ may encode superpixels; in other settings, it may correspond to a low-rank projection or simply a subset of raw features. Such simplification often makes for easier interpretation and, when $D^{\prime}\ll D$, also reduces the computational cost. See Appendix A for a summary of the notation.
@@ 2.2 Suppressor Variables
Various studies [e.g., 9, 12] demonstrated that linear models may need to assign significant non-zero weight to correlated noise variables (suppressors). Given $\mathbf{x}\in\mathbb{R}^{D}$, assume each instance is generated from a linear model of the form
$\mathbf{x}\;=\;\mathbf{A}\,\mathbf{m}\;+\;\bm{\varepsilon},\qquad\mathbf{m}=\bigl[m_{1},\dots,m_{K}\bigr]^{\!\top}\in\mathbb{R}^{K},$ (1)
where $K$ is the number of latent factors, $\mathbf{A}=[\mathbf{a}_{1},\dots,\mathbf{a}_{K}]\in\mathbb{R}^{D\times K}$ contains the activation patterns $\mathbf{a}_{k}$, and $\bm{\varepsilon}$ denotes additive noise. A linear backward model attempts to recover the latent vector $\mathbf{m}$ via
$\hat{\mathbf{m}}\;=\;\mathbf{W}^{\!\top}\mathbf{x},\qquad\mathbf{W}\in\mathbb{R}^{D\times K}.$ (2)
Consider $D=2$ and $K=1$ with $x_{1}=m+d$ and $x_{2}=d$, where the scalar $m$ is the signal of interest and $d$ is a suppressor. The signal is exactly recovered by $\hat{m}=x_{1}-x_{2}=\mathbf{w}^{\!\top}\mathbf{x}$ with $\mathbf{w}=[1,-1]^{\!\top}$. Interpreting $\mathbf{w}$ directly as feature importance, however, suggests $x_{1}$ and $x_{2}$ contribute equally, even though the second channel carries only the suppressor, which is removed entirely in $\hat{m}$ by subtraction. This issue with suppressor variables extends to common explainable AI methods used to interpret machine learning models [35, 8].
Remark.
While suppressor variables can improve prediction accuracy, they do not reflect genuine statistical dependence on the target. Such variables should be excluded from explanations, as they have no direct or indirect (causal, anti-causal, or confounded) link to the target. This differs from phenomena like the Clever Hans effect or shortcut learning, where models rely on spurious yet predictive features.
@@
Instead of using the weights of the linear backward model, Haufe et al. [12] propose to use the corresponding forward model. They show that the weight matrix $\mathbf{W}$ has a unique forward model counterpart
$\mathbf{A}\;=\;\Sigma_{\mathbf{X}}\,\mathbf{W}\,\Sigma_{\mathbf{M}}^{-1},$ (3)
where $\Sigma_{\mathbf{X}}=\operatorname{Cov}[\mathbf{x}]$ and $\Sigma_{\mathbf{M}}=\operatorname{Cov}[\hat{\mathbf{m}}$] which recovers the activation pattern in Eq.˜1. If the estimated factors are uncorrelated, $\Sigma_{\mathbf{M}}$ is diagonal and Eq.˜3 reduces to $\mathbf{A}\;\propto\;\Sigma_{\mathbf{X}}\,\mathbf{W}\;=\;\operatorname{Cov}\bigl[\mathbf{x},\,\hat{\mathbf{m}}\bigr$]. The resulting global feature importance map for the entire data set $\mathcal{D}\subset\mathcal{X}$ is
@@ 2.4 Local explainability methods
$\mathbf{s}^{\text{LIME}}(\mathbf{x}_{\star})\;=\;\mathbf{w}^{\text{LIME}}\;=\;\arg\!\min_{\mathbf{v}}\;\mathbb{E}_{\mathbf{z}^{\prime}\,\sim\,\mathbb{P}_{\mathcal{Z}}}\Bigl[\Pi_{\mathbf{x^{\prime}_{\star}}}\!\bigl(\mathbf{z}^{\prime}\bigr)\bigl(f\!\bigl(h_{\mathbf{x_{\star}}}^{-1}(\mathbf{z}^{\prime})\bigr)-\mathbf{v}^{\!\top}\mathbf{z}^{\prime}\bigr)^{2}\Bigr]\;+\;\lambda\,R(\mathbf{v}),$ (5)
where $R$ is a regularization weighted by $\lambda$, $\mathcal{Z}$ is the simplified input space, and $h_{\mathbf{x_{\star}}}^{-1}\!:\mathcal{Z}\!\rightarrow\!\mathcal{X}$ maps a simplified sample back to the original space. This notation is similar to the one by Tan et al. [31]. For image explanations, the default choices are $\Pi_{\mathbf{x^{\prime}_{\star}}}(\mathbf{z}^{\prime})=\exp\,\!(-\|\mathbf{1}-\mathbf{z}^{\prime}\|_{0}^{2}/\sigma^{2})$, $R(\mathbf{v})=\|\mathbf{v}\|_{2}^{2}$, and $h_{\mathbf{x^{\prime}_{\star}}}^{-1}(\mathbf{z}^{\prime})=\mathbf{z}=\mathbf{x_{\star}}\odot B(\mathbf{z}^{\prime})+\mathbf{r}\odot(\mathbf{1}-B(\mathbf{z}^{\prime}))$, with $\mathbf{z}^{\prime}\sim\text{Uni}(\{0,1\}^{D^{\prime}})$, $\odot$ denoting element-wise product, $r$ being a reference value, and where $B:\{0,1\}^{d}\rightarrow\{0,1\}^{D}$ copies each bit $z^{\prime}_{j}$ to every pixel in superpixel. As mentioned, LIME can assign non-zero importance to suppressor variables [35].
@@
As shown by Lundberg and Lee [17], Tan et al. [31], KernelSHAP can be written in the form of Eq.˜5 by setting $R(\mathbf{v})=0$ and $\Pi_{x^{\prime}}(z^{\prime})=(D^{\prime}-1)\big/\left[\left(D^{\prime}\,\mathrm{choose}\,\|z^{\prime}\|_{0}\right)\cdot\|z^{\prime}\|_{0}\cdot(D^{\prime}-\|z^{\prime}\|_{0})\right]$. Tan et al. [31] further demonstrate that Eq.˜5 unifies SmoothGrad [28] and the plain Gradient method [37].
@@ 3 PatternLocal
We now unify the Pattern insight that forward model parameters reveal statistically relevant features with the ability to probe highly non-linear models locally with methods like LIME. We call this PatternLocal. PatternLocal converts any linear surrogate explanation $\mathbf{w}$ produced around an instance $\mathbf{x}_{\star}$ into a data-driven pattern $\mathbf{a}$ that (1) suppresses false positives caused by suppressor variables, (2) remains faithful to the model because it builds on the very surrogate that approximates the decision boundary, (3) preserves local fidelity by weighting the surrounding data with a local kernel, and (4) remains practical in high-dimensional settings by operating in a simplified, low-dimensional space with regularized estimation.
(1) Suppressor-variable mitigation. PatternLocal mitigates suppressor effects by regressing the neighborhood of $\mathbf{x}_{\star}$ onto the local surrogate output $\tilde{y}=\mathbf{w}^{\top}\mathbf{x}$ under a Gaussian noise model. This reconstruction of $\mathbf{x}$ from $\tilde{y}$ effectively down-weights suppressor contributions.
(2) Model faithfulness. PatternLocal inherits $\mathbf{w}$ from Eq.˜5, which approximates the model $f$ locally around $\mathbf{x}_{\star}$. By regressing the simplified input $\mathbf{x}$ on the surrogate response $\tilde{y}$ rather than the true label $y$, PatternLocal ensures that explanations describe the model’s behavior, not the underlying task.
(3) Local fidelity. In non-linear models, the important features can vary across different regions of the input space. PatternLocal enforces locality by weighting all terms in its regression objective using the local kernel $\Pi_{\mathbf{x^{\prime}_{\star}}}$. This ensures that the explanation focuses on the neighborhood around $\mathbf{x}_{\star}$ and does not reflect global structure.
(4) Feasibility and dimensionality. PatternLocal assumes sufficient local samples near $\mathbf{x}_{\star}$ and operates in the simplified space defined by $\mathbf{h}_{\mathbf{x_{\star}}}$, which aligns with Eq.˜5 and reduces the dimensionality. To maintain stability in sparse or high-dimensional settings, the regression includes a regularization term $Q$.